# HAF-RM: A Hybrid Alignment Framework for Reward Model Training

### v2

@@ 5.1.1 Overall Performance
Table 2: Overall results on each dataset for accuracy, which denotes the proportion that the better response is scored higher. The best performance is highlighted in boldface and the suboptimal result is underlined.
Firstly we compare the performance of HaF with the baseline and two judging models across the five datasets. Table 2 presents the overall results. HAF has higher accuracy than the baseline in most cases, indicating that the model can more sensitively identify whether an answer is good and give a more accurate high (or low) score. At the same time, those results worse than the baseline or DPO are generally only slightly worse, indicating that our method is basically not weaker than the baseline under various circumstances.
Figure 2: Comparison of models trained with HAF/baseline/DPO methods on the mixed dataset.
Acc(%) AHP CA Helpful BS Harmless AVG rAcc
Phi-2-2.7B
AHP * $\textrm{67.40}^{{}_{(\textrm{1.00}\downarrow)}}_{{}^{(\textrm{30.30}\uparrow)}}$ $\textrm{67.60}^{{}_{(\textrm{3.40}\uparrow)}}_{{}^{(\textrm{17.10}\uparrow)}}$ $\textrm{39.80}^{{}_{(\textrm{0.20}\uparrow)}}_{{}^{(\textrm{14.80}\downarrow)}}$ $\textrm{41.90}^{{}_{(\textrm{5.40}\downarrow)\phantom{1}}}_{{}^{(\textrm{9.30}\downarrow)\phantom{1}}}$ $\textrm{54.18}^{{}_{(\textrm{0.70}\downarrow)\phantom{1}}}_{{}^{(\textrm{5.83}\uparrow)\phantom{1}}}$ $\textrm{67.50}^{{}_{(\textrm{1.20}\uparrow)}}_{{}^{(\textrm{23.70}\uparrow)}}$
CA $\textrm{60.20}^{{}_{(\textrm{0.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{8.20}\uparrow)\phantom{1}}}$ $\textrm{64.70}^{{}_{(\textrm{3.20}\downarrow)}}_{{}^{(\textrm{15.00}\uparrow)}}$ $\textrm{37.60}^{{}_{(\textrm{0.80}\uparrow)}}_{{}^{(\textrm{13.00}\downarrow)}}$ $\textrm{42.10}^{{}_{(\textrm{5.60}\uparrow)\phantom{1}}}_{{}^{(\textrm{9.40}\downarrow)\phantom{1}}}$ $\textrm{51.15}^{{}_{(\textrm{0.92}\uparrow)\phantom{1}}}_{{}^{(\textrm{0.20}\uparrow)\phantom{1}}}$ $\textrm{62.45}^{{}_{(\textrm{1.35}\downarrow)}}_{{}^{(\textrm{11.60}\uparrow)}}$
Helpful $\textrm{60.20}^{{}_{(\textrm{2.90}\uparrow)\phantom{1}}}_{{}^{(\textrm{6.90}\uparrow)\phantom{1}}}$ $\textrm{72.00}^{{}_{(\textrm{1.10}\downarrow)}}_{{}^{(\textrm{32.70}\uparrow)}}$ $\textrm{36.20}^{{}_{(\textrm{1.40}\downarrow)}}_{{}^{(\textrm{10.30}\downarrow)}}$ $\textrm{38.50}^{{}_{(\textrm{6.90}\downarrow)\phantom{1}}}_{{}^{(\textrm{0.30}\downarrow)\phantom{1}}}$ $\textrm{51.73}^{{}_{(\textrm{1.62}\downarrow)\phantom{1}}}_{{}^{(\textrm{7.25}\uparrow)\phantom{1}}}$ $\textrm{66.10}^{{}_{(\textrm{0.90}\uparrow)}}_{{}^{(\textrm{19.80}\uparrow)}}$
BS $\textrm{47.90}^{{}_{(\textrm{0.20}\downarrow)\phantom{1}}}_{{}^{(\textrm{1.20}\downarrow)\phantom{1}}}$ $\textrm{41.00}^{{}_{(\textrm{2.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{9.20}\uparrow)\phantom{1}}}$ $\textrm{35.70}^{{}_{(\textrm{1.40}\downarrow)\phantom{1}}}_{{}^{(\textrm{9.30}\downarrow)\phantom{1}}}$ $\textrm{70.60}^{{}_{(\textrm{5.60}\uparrow)\phantom{1}}}_{{}^{(\textrm{4.60}\uparrow)\phantom{1}}}$ $\textrm{48.80}^{{}_{(\textrm{1.62}\uparrow)\phantom{1}}}_{{}^{(\textrm{0.82}\uparrow)\phantom{1}}}$ $\textrm{70.60}^{{}_{(\textrm{5.60}\uparrow)\phantom{1}}}_{{}^{(\textrm{4.60}\uparrow)\phantom{1}}}$
Harmless $\textrm{43.80}^{{}_{(\textrm{1.30}\uparrow)\phantom{1}}}_{{}^{(\textrm{6.20}\downarrow)\phantom{1}}}$ $\textrm{29.40}^{{}_{(\textrm{0.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{5.70}\downarrow)\phantom{1}}}$ $\textrm{32.60}^{{}_{(\textrm{0.80}\uparrow)\phantom{1}}}_{{}^{(\textrm{9.10}\downarrow)\phantom{1}}}$ $\textrm{76.90}^{{}_{(\textrm{1.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{8.60}\uparrow)\phantom{1}}}$ $\textrm{45.67}^{{}_{(\textrm{1.02}\uparrow)\phantom{1}}}_{{}^{(\textrm{3.10}\downarrow)\phantom{1}}}$ $\textrm{76.90}^{{}_{(\textrm{1.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{8.60}\uparrow)\phantom{1}}}$
Mistral-7B-Instruct
AHP * $\textrm{75.50}^{{}_{(\textrm{6.20}\uparrow)}}_{{}^{(\textrm{17.90}\uparrow)}}$ $\textrm{68.90}^{{}_{(\textrm{10.60}\uparrow)}}_{{}^{(\textrm{7.60}\uparrow)}}$ $\textrm{55.70}^{{}_{(\textrm{7.20}\uparrow)\phantom{1}}}_{{}^{(\textrm{5.90}\uparrow)\phantom{1}}}$ $\textrm{48.00}^{{}_{(\textrm{1.20}\uparrow)\phantom{1}}}_{{}^{(\textrm{1.40}\downarrow)\phantom{1}}}$ $\textrm{62.02}^{{}_{(\textrm{6.30}\uparrow)\phantom{1}}}_{{}^{(\textrm{7.50}\uparrow)\phantom{1}}}$ $\textrm{72.20}^{{}_{(\textrm{8.40}\uparrow)}}_{{}^{(\textrm{12.75}\uparrow)}}$
CA $\textrm{60.80}^{{}_{(\textrm{0.20}\downarrow)\phantom{1}}}_{{}^{(\textrm{6.80}\uparrow)\phantom{1}}}$ $\textrm{65.80}^{{}_{(\textrm{1.20}\downarrow)}}_{{}^{(\textrm{12.50}\uparrow)}}$ $\textrm{38.50}^{{}_{(\textrm{6.60}\downarrow)\phantom{1}}}_{{}^{(\textrm{3.00}\downarrow)\phantom{1}}}$ $\textrm{36.80}^{{}_{(\textrm{4.00}\downarrow)\phantom{1}}}_{{}^{(\textrm{6.80}\downarrow)\phantom{1}}}$ $\textrm{50.47}^{{}_{(\textrm{3.00}\downarrow)\phantom{1}}}_{{}^{(\textrm{2.37}\uparrow)\phantom{1}}}$ $\textrm{63.30}^{{}_{(\textrm{0.70}\downarrow)\phantom{1}}}_{{}^{(\textrm{9.65}\uparrow)\phantom{1}}}$
Helpful $\textrm{60.90}^{{}_{(\textrm{1.00}\downarrow)\phantom{1}}}_{{}^{(\textrm{8.20}\uparrow)\phantom{1}}}$ $\textrm{73.90}^{{}_{(\textrm{0.60}\uparrow)}}_{{}^{(\textrm{20.30}\uparrow)}}$ $\textrm{36.00}^{{}_{(\textrm{9.30}\downarrow)\phantom{1}}}_{{}^{(\textrm{5.40}\downarrow)\phantom{1}}}$ $\textrm{37.50}^{{}_{(\textrm{0.00})\phantom{1}}}_{{}^{(\textrm{2.10}\uparrow)\phantom{1}}}$ $\textrm{52.08}^{{}_{(\textrm{2.42}\downarrow)\phantom{1}}}_{{}^{(\textrm{6.30}\uparrow)\phantom{1}}}$ $\textrm{67.40}^{{}_{(\textrm{0.20}\downarrow)}}_{{}^{(\textrm{14.25}\uparrow)}}$
BS $\textrm{52.90}^{{}_{(\textrm{4.00}\uparrow)\phantom{1}}}_{{}^{(\textrm{1.10}\downarrow)\phantom{1}}}$ $\textrm{55.20}^{{}_{(\textrm{9.50}\uparrow)}}_{{}^{(\textrm{12.30}\uparrow)}}$ $\textrm{43.80}^{{}_{(\textrm{3.40}\downarrow)\phantom{1}}}_{{}^{(\textrm{8.10}\downarrow)\phantom{1}}}$ $\textrm{71.90}^{{}_{(\textrm{1.40}\uparrow)\phantom{1}}}_{{}^{(\textrm{3.00}\uparrow)\phantom{1}}}$ $\textrm{55.95}^{{}_{(\textrm{2.87}\uparrow)\phantom{1}}}_{{}^{(\textrm{1.52}\uparrow)\phantom{1}}}$ $\textrm{71.90}^{{}_{(\textrm{1.40}\uparrow)\phantom{1}}}_{{}^{(\textrm{3.00}\uparrow)\phantom{1}}}$
Harmless $\textrm{46.50}^{{}_{(\textrm{1.00}\uparrow)\phantom{1}}}_{{}^{(\textrm{0.40}\downarrow)\phantom{1}}}$ $\textrm{38.30}^{{}_{(\textrm{4.60}\uparrow)}}_{{}^{(\textrm{10.50}\downarrow)}}$ $\textrm{32.40}^{{}_{(\textrm{0.50}\uparrow)\phantom{1}}}_{{}^{(\textrm{2.10}\downarrow)\phantom{1}}}$ $\textrm{76.70}^{{}_{(\textrm{2.40}\uparrow)\phantom{1}}}_{{}^{(\textrm{5.70}\uparrow)\phantom{1}}}$ $\textrm{48.48}^{{}_{(\textrm{2.13}\uparrow)\phantom{1}}}_{{}^{(\textrm{1.82}\downarrow)\phantom{1}}}$ $\textrm{76.70}^{{}_{(\textrm{2.40}\uparrow)\phantom{1}}}_{{}^{(\textrm{5.70}\uparrow)\phantom{1}}}$
@@
Table 3: Results for the OOD experiment. The results in the same row are derived from the same backbone and the same training dataset, while the columns represent different test datasets. The displayed accuracies are for HaF , with superscripts and subscripts indicating the performance differences relative to the baseline and DPO, respectively. $\uparrow$ denotes an improvement with HaF , whereas $\downarrow$ indicates a decline. rAcc is the average accuracy among grey blocks.
5.1.2 Mixed Data
@@
For the mixed data setting, we construct two datasets by sampling and combining examples from multiple sources: Anthropic-HH (Anthropic Helpful + Anthropic Harmless) and Mixed (evenly sampled from each of the five datasets in our corpus). As shown in Figure 2, our proposed hybrid alignment framework achieves the best generalization performance across all reward models when evaluated on these mixed data distributions. This suggests our approach can better learn the diversity present in the combined datasets for generalization.
5.1.3 OOD Data
Data within the same dataset often exhibits certain distributional similarities due to similar or even identical data cleaning and processing methods. To simulate a distribution shift in real-world application, we also evaluate generalization to entirely held-out OOD datasets. Specifically, we train models on one dataset and evaluate on the remaining four. Although different datasets have distinct distributions, their main preferences can be generalized as “overall better” (AHP, CA and Helpful) and “safer” (BS and Harmless), which we use rAcc (“r” stands for “relevant”) to represent model’s generalization ability within similar preferences.
@@ 5.1.3 OOD Data
The results are detailed in Table 3. We can tell from the table that the rAcc of HaF is basically higher than that of both Baseline and DPO, indicating HaF possesses a strong ability to learn preferences and effectively generalize them to similar preference distributions, despite great differences in language style and topic. Touvron et al. (2023) noted that RLHF involves distributional shifts in the policy model during training, necessitating iterative training for the reward model. The robustness of HaF against such distributional shifts could potentially be a key factor in alleviating this problem.
Comparing Table 3 and Table 2, we can observe that models trained using CA or Helpful datasets outperform those directly trained on AHP dataset when the test set is AHP. This suggests two things: firstly, there is a certain similarity in preferences across the three datasets, and secondly, the amount of AHP data may be insufficient to support the complete training of the reward model, as shown in Table 1. Consequently, the model’s preference learning is incomplete, which results in low test outcomes for AHP in Table 2.
On this observation, it can be noted that the HaF-mistral model, when fully trained using CA or Helpful datasets, performs worse on BS and Harmless compared to Baseline. However, when insufficiently trained using AHP dataset, its test results are better than the baseline. This might indicate that during the training process of the reward model, the learning of reward mapping precedes the learning of preferences. When the model is not fully trained, HAF’s advantage in learning speed enables it to outperform. Yet, once fully trained, the baseline’s weaker preference learning ability might allow it to exhibit some degree of cross-preference generalization. This hypothesis requires further validation in future work.
@@
One easily overlooked result is that nearly all the test outcomes of the DPO model converge to approximately 50% in a highly exaggerated manner, indicating a complete loss of modeling capability for out-of-distribution data. This issue is likely related to its inherent nature as a language model: the generation process of language models exhibits strong stylistic tendencies, which, in turn, leads to a significantly higher preference for responses that align with its style (as reflected in the generation probabilities and the implicit reward values of the DPO model). Consequently, when the response distribution deviates from its stylistic norms (e.g., responses that are too short or too long, or use different vocabulary), the output probabilities become highly inaccurate. This indicates that the DPO model is not suitable for use as a conventional reward model.
@@ 5.2 Extrinsic Evaluation on Downstream Task
In assessing the practical applicability of reward models, intrinsic performances alone provide an incomplete picture of their efficacy. To comprehensively evaluate their utility in real-world applications, it is essential to examine how these models perform in downstream tasks that simulate practical scenarios.
This section aims to investigate the robustness and effectiveness of HaF model in such scenarios. Specifically, we explore its performance in two distinct downstream tasks: best-of-N sampling as a training-free response generation strategy (Stiennon et al., 2020; Gao et al., 2023; Jinnai et al., 2024), and RLHF as a training-dependent aligning methods.

### v3

@@ 4.1.1 Overall Performance
Table 2: Overall results (accuracy) for each dataset, by calculating the proportion that the better response is scored higher. The best performance is highlighted in boldface and the suboptimal result is underlined.
Table 2 presents the overall results of our HaF compared to two basic approaches across five datasets. We observe that DPO and the baseline method show similar performance on average but there is significant variability in individual comparisons. This suggests that the two methods focus on different features when learning preferences. In contrast, HAF consistently outperforms both, indicating its ability to effectively integrate features from both approaches to better learn preferences.
Specifically, Mistral-base performs poorly on the Helpful, CA, and AHP datasets because these datasets require preferences related to the quality of responses. Since the base model has not undergone instruction tuning, it lacks the representation of relevant features, making it difficult to accurately judge response quality. In contrast, the extensively trained base model is capable of distinguishing between benign and harmful content, allowing it to perform comparably to Mistral-Instruct on the safety-related BS and Harmless datasets. Nevertheless, HaF demonstrates promising results even for these challenging preferences.
Notably, DPO achieves the highest performance on BS across all three models, which is probably caused by DPO’s “concentrated” data-fitting manner (Azar et al., 2023). This is evident from the much lower variance in token-level perplexity for good and bad responses in the BS dataset compared to other datasets, indicating a more concentrated distribution respectively of these two subsets (refer to Appendix E.2 for detailed illustration). By integrating DPO loss, our HaF partially captures this “concentrated” data-fitting characteristics, leading to a more nuanced improvement on BS compared to the baseline methods. However, DPO’s concentrated data-fitting may potential lead to over-fitting issues, whereas HaF and the baseline demonstrate better generalization ability, which we will elaborate on in the following experiments.
Figure 3: The performance differences of HaF / baseline / DPO under mixed preference training, with light shading indicating the upper bound performance of individually trained reward models on each dataset.
Acc(%) AHPC CAC HelpfulC BSS HarmlessS Avg.
internal
Phi-2 $\textrm{67.50}^{{}_{(\textrm{1.20}\uparrow)}}_{{}^{(\textrm{23.70}\uparrow)}}$ $\textrm{62.45}^{{}_{(\textrm{1.35}\downarrow)}}_{{}^{(\textrm{11.60}\uparrow)}}$ $\textrm{66.10}^{{}_{(\textrm{0.90}\uparrow)}}_{{}^{(\textrm{19.80}\uparrow)}}$ $\textrm{70.60}^{{}_{(\textrm{5.60}\uparrow)}}_{{}^{(\textrm{4.60}\uparrow)}}$ $\textrm{76.90}^{{}_{(\textrm{1.50}\uparrow)}}_{{}^{(\textrm{8.60}\uparrow)}}$ $\textrm{68.71}^{{}_{(\textrm{1.57}\uparrow)}}_{{}^{(\textrm{13.66}\uparrow)}}$
Mistral-base $\textrm{59.65}^{{}_{(\textrm{4.90}\uparrow)}}_{{}^{(\textrm{14.55}\uparrow)}}$ $\textrm{56.35}^{{}_{(\textrm{2.15}\downarrow)}}_{{}^{(\textrm{6.00}\uparrow)}}$ $\textrm{62.40}^{{}_{(\textrm{0.85}\downarrow)}}_{{}^{(\textrm{12.85}\uparrow)}}$ $\textrm{69.60}^{{}_{(\textrm{0.50}\uparrow)}}_{{}^{(\textrm{3.30}\uparrow)}}$ $\textrm{75.30}^{{}_{(\textrm{1.90}\uparrow)}}_{{}^{(\textrm{5.80}\uparrow)}}$ $\textrm{64.66}^{{}_{(\textrm{0.86}\uparrow)}}_{{}^{(\textrm{8.50}\uparrow)}}$
Mistral $\textrm{72.20}^{{}_{(\textrm{8.40}\uparrow)}}_{{}^{(\textrm{12.75}\uparrow)}}$ $\textrm{63.30}^{{}_{(\textrm{0.70}\downarrow)}}_{{}^{(\textrm{9.65}\uparrow)}}$ $\textrm{67.40}^{{}_{(\textrm{0.20}\downarrow)}}_{{}^{(\textrm{14.25}\uparrow)}}$ $\textrm{71.90}^{{}_{(\textrm{1.40}\uparrow)}}_{{}^{(\textrm{3.00}\uparrow)}}$ $\textrm{76.70}^{{}_{(\textrm{2.40}\uparrow)}}_{{}^{(\textrm{5.70}\uparrow)}}$ $\textrm{70.30}^{{}_{(\textrm{2.26}\uparrow)}}_{{}^{(\textrm{9.07}\uparrow)}}$
external
Phi-2 $\textrm{85.14}^{{}_{(\textrm{1.36}\uparrow)}}_{{}^{(\textrm{65.88}\uparrow)}}$ $\textrm{95.27}^{{}_{(\textrm{0.34}\uparrow)}}_{{}^{(\textrm{19.59}\uparrow)}}$ $\textrm{89.86}^{{}_{(\textrm{6.08}\uparrow)}}_{{}^{(\textrm{74.66}\uparrow)}}$ $\textrm{66.30}^{{}_{(\textrm{0.95}\uparrow)}}_{{}^{(\textrm{2.04}\uparrow)}}$ $\textrm{66.44}^{{}_{(\textrm{0.38}\downarrow)}}_{{}^{(\textrm{4.62}\uparrow)}}$ $\textrm{80.60}^{{}_{(\textrm{8.35}\uparrow)}}_{{}^{(\textrm{33.36}\uparrow)}}$
Mistral-base $\textrm{79.66}^{{}_{(\textrm{20.34}\uparrow)}}_{{}^{(\textrm{64.14}\uparrow)}}$ $\textrm{93.79}^{{}_{(\textrm{21.03}\uparrow)}}_{{}^{(\textrm{33.45}\uparrow)}}$ $\textrm{81.38}^{{}_{(\textrm{6.90}\downarrow)}}_{{}^{(\textrm{67.24}\uparrow)}}$ $\textrm{70.40}^{{}_{(\textrm{3.27}\uparrow)}}_{{}^{(\textrm{8.73}\uparrow)}}$ $\textrm{63.30}^{{}_{(\textrm{3.82}\downarrow)}}_{{}^{(\textrm{4.91}\uparrow)}}$ $\textrm{77.70}^{{}_{(\textrm{6.79}\uparrow)}}_{{}^{(\textrm{35.69}\uparrow)}}$
Mistral $\textrm{91.55}^{{}_{(\textrm{32.77}\uparrow)}}_{{}^{(\textrm{53.37}\uparrow)}}$ $\textrm{91.89}^{{}_{(\textrm{3.04}\uparrow)}}_{{}^{(\textrm{16.21}\uparrow)}}$ $\textrm{82.43}^{{}_{(\textrm{1.69}\uparrow)}}_{{}^{(\textrm{63.51}\uparrow)}}$ $\textrm{70.52}^{{}_{(\textrm{1.22}\downarrow)}}_{{}^{(\textrm{4.08}\uparrow)}}$ $\textrm{73.37}^{{}_{(\textrm{2.72}\uparrow)}}_{{}^{(\textrm{5.17}\uparrow)}}$ $\textrm{81.95}^{{}_{(\textrm{7.80}\uparrow)}}_{{}^{(\textrm{28.47}\uparrow)}}$
@@
Table 3: Results for out-of-distribution data. Subscripts $C$ and $S$ denote the subjects of training sets, where $C$ represents Chat and $S$ represents Safety. “internal” refers to testing results among datasets sharing the same subject category, while “external” refers to testing results on RewardBench. The displayed accuracies are for HaF , with superscripts and subscripts indicating the performance differences relative to the baseline and DPO, respectively. $\uparrow$ denotes an improvement with HaF , while $\downarrow$ signifies a decline.
4.1.2 Evaluation on Mixed Data
@@
To illustrate HaF ’s effectiveness in training reward models on mixed data, we construct a dataset by evenly sampling and combining examples from all five datasets. As shown in Figure 3, our proposed hybrid alignment framework achieves the best overall performance across all reward models when evaluated on the mixed data distribution. This suggests that HaF is more effective at learning the diversity within the combined datasets.
Specifically, compared to the individual results on corresponding datasets in Table 2 (shown as lightly shaded bars in Figure 3), we observe that both the baseline method and HaF replicate their performance in learning individual preferences better than DPO when applied to mixed preference learning. Notably, DPO’s performance drops significantly on the CA and Helpful datasets, suggesting that DPO tends to fit the most prominent features of the overall data distribution. This also aligns with the finding of Chen et al. (2024) that DPO would optimize the margins of correct data rather than the wrong ones.
4.1.3 Transferability to OOD Data
We further evaluate the generalizability of our framework to entirely held-out out-of-distribution (OOD) datasets to simulate distribution shifts in real-world applications. Specifically, the five datasets are grouped into two categories: “Safety” (BS, Harmless) and “Chat” (AHP, CA, Helpful). We train the model on one dataset and evaluate its performance within the same category. The evaluation data comes from two sources, including the “internal” source referring to different datasets within the same category, and an “external” source, consisting of test data on related topics from RewardBench.
@@ 4.1.3 Transferability to OOD Data
As shown in Table 3, HaF achieves a higher internal accuracy compared to both Baseline and DPO, demonstrating HaF ’s strong ability to learn preferences and effectively generalize to similar preference distributions, even with notable differences in language style and topic. As Touvron et al. (2023) noted, RLHF causes distributional shifts in the policy model during training, often requiring iterative training of the reward model. HaF ’s robustness against these distributional shifts could potentially be a key factor in mitigating this issue.
@@
It is important to note that nearly all of DPO’s test outcomes converge around 50%, indicating a complete loss of modeling capability for OOD data. This likely stems from DPO’s inherent nature as a language model, where the generation process exhibits strong stylistic biases, favoring responses that align with its style (as reflected in generation probabilities and implicit reward values). When response distribution deviates from these stylistic norms (e.g., responses that are too short, too long, or use different vocabulary), DPO’s output probabilities become highly inaccurate, rendering it unsuitable as a conventional reward model.
From these three experiments, we conclude that DPO learns features significantly different from those learned by the baseline method. In contrast, HaF inherits both the baseline method’s generalization ability and DPO’s stronger fitting capability.
@@ 4.2 Extrinsic Evaluation on Downstream Task
Intrinsic performance metrics offer only a partial view of a reward model’s efficacy. To comprehensively assess their practical applicability in real-world scenarios, it is crucial to evaluate how these models perform in downstream tasks that closely simulate practical applications.
In this section, we evaluate the robustness and effectiveness of HaF in such scenarios. Specifically, we explore its performance in two distinct downstream tasks: best-of-N sampling, a training-free response generation strategy (Stiennon et al., 2020; Gao et al., 2023; Jinnai et al., 2024), and RLHF, a training-dependent alignment method.